A construction company is purchasing a new Tractor for over the road use. The IRS classifies this as 3-year property. The truck costs $471000.

Determine the depreciation allowance for each year using DDB method.
Year 1 $  
Year 2 $  
Year 3 $  
Year 4 $  
Answer #1

P = 471000, life = 3 yrs

d = 2/n = 2/3

depreciation in year 1 = 2/3 * 471000 = 314000

Book value at the end of yr 1 = 471000 - 314000 = 157000

Depreciation in yr 2 = 2/3 * 157000 = 104666.67

Book value at the end of yr 2 = 157000 - 104666.67 = 52333.33

Depreciation in yr 3 = 2/3 * 52333.33 = 34888.89

Book value at the end of yr 3 = 52333.33 - 34888.89 = 17444.44

Depreciation in yr 4 = 2/3 * 17444.44 = 11629.63